OpenStudy (anonymous):

please help solve. find the center and radius of the circle. x^2 + y^2 - 10x - 16y + 80 = 0 please show steps

ganeshie8 (ganeshie8):

put x terms together put y terms together complete teh square for each

ganeshie8 (ganeshie8):

x^2 + y^2 - 10x - 16y + 80 = 0 x^2 - 10x + y^2 - 16y = -80

ganeshie8 (ganeshie8):

familiear with completing the square ?

OpenStudy (anonymous):

im not that great at it

ganeshie8 (ganeshie8):

give it a try

ganeshie8 (ganeshie8):

x^2 - 10x take half of 10, you get 5. square it and add both sides

ganeshie8 (ganeshie8):

do the same for y^2 - 16y

ganeshie8 (ganeshie8):

x^2 + y^2 - 10x - 16y + 80 = 0 x^2 - 10x + y^2 - 16y = -80 x^2 - 10x + `5^2` + y^2 - 16y + `8^2` = -80 + `5^2` + `8^2` (x-5)^2 + (y-8)^2 = -80 + 25 + 64 (x-5)^2 + (y-8)^2 = 9

OpenStudy (anonymous):

so its not x=4+- 10^2

ganeshie8 (ganeshie8):

thats it! see if that makes more or less sense..

ganeshie8 (ganeshie8):

so the equation of circle in center radius form is : (x-5)^2 + (y-8)^2 = 3^2 center = (5, 8) radius = 3

ganeshie8 (ganeshie8):

let me knw if something doesn't make sense

OpenStudy (anonymous):

how you get 5 for completing the square?

ganeshie8 (ganeshie8):

thats a very good question

ganeshie8 (ganeshie8):

x^2 + y^2 - 10x - 16y + 80 = 0 x^2 - 10x + y^2 - 16y = -80 x^2 - 10x + `5^2` + y^2 - 16y + `8^2` = -80 + `5^2` + `8^2`

ganeshie8 (ganeshie8):

you're fine till this step ?

OpenStudy (anonymous):

ya thats where i get stuck i can get everything past that part

ganeshie8 (ganeshie8):

you're fine with 3rd line, right ?

ganeshie8 (ganeshie8):

x^2 + y^2 - 10x - 16y + 80 = 0 x^2 - 10x + y^2 - 16y = -80 x^2 - 10x + `5^2` + y^2 - 16y + `8^2` = -80 + `5^2` + `8^2`

ganeshie8 (ganeshie8):

all above lines make sense ?

OpenStudy (anonymous):

i dont get the completing the square part how you get the 5 and 8

ganeshie8 (ganeshie8):

x^2 - 10x 5 is half of x coefficient

ganeshie8 (ganeshie8):

y^2 - 16y 8 is half of y coefficient

ganeshie8 (ganeshie8):

its just a process, memorize it for now

OpenStudy (anonymous):

so its just half of y?

ganeshie8 (ganeshie8):

its half of y coefficient

OpenStudy (anonymous):

k i think i got it... the equation made it look like there was more to it